# Can Language Models Replace Programmers? RepoCod Says ‘Not Yet’

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable ability in code generation with more than 90 pass@1 in solving Python coding problems in HumanEval and MBPP. Such high accuracy leads to the question: can LLMs replace human programmers? Existing manual crafted, simple, or single-line code generation benchmarks cannot answer this question due to their gap with real-world software development. To answer this question, we propose RepoCod, a code generation benchmark with 980 problems collected from 11 popular real-world projects, with more than 58% of them requiring file-level or repository-level context information. In addition, RepoCod has the longest average canonical solution length (331.6 tokens) and the highest average cyclomatic complexity (9.00) compared to existing benchmarks. In our evaluations on ten LLMs, none of the models can achieve more than 30 pass@1 on RepoCod, disclosing the necessity of building stronger LLMs that can help developers in real-world software development.
@@ 1 Introduction
The contributions of our work are summarized as follows:
  • A challenging code generation dataset incorporates developer-defined test cases as evaluation metrics, reflecting real-world complexity.
@@
    • There are 257 (26.22%) of instances in RepoCod that require repository-level contexts.
@@
    • LLMs are bad at generating functions with high content complexity (up to 20 pass@1 for problems requiring repository-level context).

### v2

@@ Abstract
Large language models (LLMs) have achieved high accuracy, i.e., more than 90 pass@1, in solving Python coding problems in HumanEval and MBPP. Thus, a natural question is, whether LLMs achieve comparable code completion performance compared to human developers? Unfortunately, one cannot answer this question using existing manual crafted or simple (e.g., single-line) code generation benchmarks, since such tasks fail to represent real-world software development tasks. In addition, existing benchmarks often use poor code correctness metrics, providing misleading conclusions.
To address these challenges, we create RepoCod, a code generation benchmark with 980 problems collected from 11 popular real-world projects, with more than 58% of them requiring file-level or repository-level context information. In addition, RepoCod has the longest average canonical solution length (331.6 tokens) and the highest average cyclomatic complexity (9.00) compared to existing benchmarks. Each task in RepoCod includes 313.5 developer-written test cases on average for better correctness evaluation. In our evaluations of ten LLMs, none of the models achieve more than 30 pass@1 on RepoCod, indicating the necessity of building stronger LLMs that can help developers in real-world software development. RepoCod is available at https://github.com/lt-asset/REPOCOD
@@ 1 Introduction
This paper makes the following contributions:
  • A challenging code generation dataset, consisting of real-world complex coding tasks, with developer-created tests for better correctness evaluation.
@@
    • Over one quarter (257, 26.22%) of instances in RepoCod require repository-level contexts.
@@
    • LLMs are ineffective at generating functions with high content complexity (up to 20 pass@1 for tasks requiring repository-level context).